OpenStudy (anonymous):

In the 2007 season, Magglio Ordonez of the Detroit tigers had 216 hits. His batting average was 0.363, the highest in major league baseball for that season. This means that of the total number of bats, 36.3% were hits. How many bats did he have?

OpenStudy (anonymous):

216 / 36.3% or 216 / .363 = 595 check 595 * 36.3% = 216

OpenStudy (anonymous):

you can set it up like this generally: \[216 / x = 36.3/100\] and solve for x

OpenStudy (anonymous):

Oh ok thanks man, i normally know how to do problems like this except i completely brain farted when i got 595 as a decimal. thanks buddy appreciate it